# Evaluating crop rotations around the world using satellite imagery and causal machine learning

### v1

@@ 4.5 Analyzing heterogeneity of rotation benefits with weather
The above regression includes a linear time trend to control for gradually changing factors besides weather that may affect rotation benefits (e.g., soil quality, management practices, crop variety choices, etc.), and to reduce how much such factors can confound our estimates of how rotation benefits vary with weather. The regression was fit on the subsample of points where the two year sequence was either AB or BB and the estimated propensity score was between and . Instead of implementing this regression directly, we implemented the regression using the best_linear_projection function in the grf package [62] which provided standard errors and -values that appropriately accounted for the fact that the same sample was being used to estimate the response variable () and to estimate the regression coefficients themselves. The best_linear_projection function also appropriately accounted for the large errors (associated with nonparametric estimation) in the estimates of when computing standard errors and -values of the regression coefficients in (3). In addition, we used the default setting where standard errors and -values for the regression coefficients were based on heteroskedasticity-robust (HC3) estimation, although they did not account for potential spatial correlations (see Section A.5). We refer readers to the documentation of the grf package [63] and [70] for more details on how this regression was fit using an alternative regression that appropriately accounts for errors in estimation of and to [71] for more details on how HC3 standard errors are calculated in regression settings.
@@
We repeated this process for each precrop A, outcome crop B, and country listed in Table A8. The rescaled heterogeneity coefficients can be roughly interpreted as the amount by which the estimated benefit of rotation on our normalized yield proxy are expected to change as the growing season precipitation (or temperature) increases from the 25th percentile to the 75th percentile in the regions where both the rotation AB and the no rotation control BB are commonly observed.
@@
We remark that our approaches do not depend upon linearity assumptions. In spite of our use of a linear model in Equation (3) to study heterogeneity with weather, we do not use linear approaches for the initial estimation of or the treatment effects (e.g., [72]) to ensure that our results on the precrop and diversification effects (Figures 3 and 5; Table A8 (Columns 6 and 7)) do not lean on any linearity assumptions. Moreover, the population regression coefficients giving the best fit to model (3) are still well defined estimands of interest regardless of whether or not is linear in temperature and precipitation. Regardless of whether is linear, the estimands are the well-defined solutions to
@@
and moreover, such estimands still provide useful summary statistics about how varies with precipitation and temperature. We use linear approaches to study heterogeneity in in part because it allows valid -value and confidence interval construction. In particular, the best_linear_projection function we use returns confidence intervals that account for the fact that is not known precisely and was estimated from the same sample, whereas existing software for nonlinear approaches (e.g., partial dependence plots, accumulated local effect plots, Shapley values) do not return such confidence intervals.
4.6 Assessment of diverse crop rotations
@@
For each two year crop sequence AB with A and B being distinct, we also assessed whether additional rotational diversity benefits crop yields. To do this, we considered samples with a three year cropping sequence of BAB as simple rotation control units and set on such samples. All samples with a three year cropping sequence of DAB, where D is any croptype that is distinct from A and B were considered as treated units with (we remark that samples of the form DAB where D was fallow, pasture or some other non-crop land classification were not included as treated units and were removed from this analysis (see Appendix G)). With a binary treatment variable distinguishing simple and diverse rotations and an outcome variable giving the GCVI-based proxy (defined in Section 4.2) for B’s yield in the third year of the crop sequence, we ran a causal forest and estimated the ATO using the same software and weather controls described in Section 4.4. The ATO estimates provided estimates of the benefit of additional rotational diversity compared to a less diverse BAB cropping sequence. While there are many possible diverse rotations with different choices of crops to be grown prior to the AB sequence that can be studied, our ATO estimate gave a weighted average of the diversification benefit across all the crops distinct from A and B that are grown before the AB sequence (with weights proportional to the prevalence in our sample).
@@
4.7 Converting results to the crop yield scale
Our estimated effects of rotation and diversification described in Sections 4.4 and 4.6 and our rescaled heterogeneity coefficients described in Section 4.5 each used normalized peak GCVI as the outcome variable. To improve the interpretability of our results, we converted these estimated effects and heterogeneity coefficients from a unitless GCVI scale to the crop yield scale by using a post-hoc, linear calibration with subnational-level crop yield data.
In particular, to calibrate our results to the yield scale in the United States, for each crop type , county , and year we computed the average normalized peak GCVI, which we call using the sample described in Section 4.1. We then extracted annual county-level crop yield data from the National Agricultural Statistics Service of the United States Department of Agriculture [73], which we call . Then for each crop type , we fit the following weighted linear regression to estimate calibration coefficients
@@ 4.7 Converting results to the crop yield scale
Due to lack of availability of sufficiently high resolution subnational-level yield data in China, we were unable to deploy the above approach in each country. Instead, we used the scaling factors of learned in the US to convert our results to the yield scale (in units of percent of average yield) in each country for each outcome crop . We used subnational-level yield data in Canada and France to validate this approach and found that subnational-level yield predictions in Canada and France that were based on coefficients from a calibration fit in the US were only slightly less accurate than those based on a calibration fit with local (i.e., from Canada or France) subnational-level yield data (see Appendix H for details).
@@
We remark that we used weighted regression in our calibration to avoid giving too much weight to counties and years that were underrepresented in our US data sample. We also note that we rescale our estimates and standard errors by a factor of rather than by for two reasons. First, the factor that we use converts our estimated average effects (or estimated average changes in effects with weather) to be in units of percent of average yield rather than units of tons per hectare. The former units have been used in recent work studying the impacts of tillage on crop yield [43] and are more interpretable in comparisons across outcome crops. Second, dividing by leads to a more robust calibration in the sense that the calibration learned in the United States using and can more reliably be used in other countries than can a calibration using just and (Appendix H). Finally, we remark that interested readers can convert our results presented in units of percent of average yield to units of tons per hectare using the average yield estimates for our sample presented in Table A9.
4.8 Net rotation benefit versus direct rotation benefit terminology
@@
Throughout the text, we present estimates of the net effects of crop rotation on yield and how those net effects vary with weather. The net effects consist of the direct effect of rotation on crop yield as well as the presumably negative indirect effect of crop rotation on yield that is mediated by changes in chemical input use. Throughout the text, in cases where it is important to distinguish between the net effects on yields that we estimate and the direct effects on yields (i.e., the effect of crop rotation on yield at fixed chemical input levels) that are often studied in field experiments, we refer to the latter as “direct rotation benefits”. The net rotation (or diversification) benefits we estimate are essentially the benefit of crop rotation (or diversification) on yield when the treated units (i.e., the rotated (or diversified) units) are defined to also use their actual reduction in chemical inputs compared to the control units (i.e., the nonrotated (or non-diversified) units). The rotation (and diversification) benefits we study are causal quantities of interest, but we note that they only capture the yield benefits of rotation and do not capture additional benefits associated with reduced chemical input costs.
Data and code availability
Acknowledgments
This work was supported by a Stanford Interdisciplinary Graduate Fellowship and the NASA Harvest and ACRES Consortia (NASA Applied Sciences Grant No. 80NSSC23M0032 (sub-award 125062-Z6521205) and 80NSSC23M0034 (sub-award 124245-Z6512205)). The authors wish to thank Kevin Guo and Stefan Wager for helpful discussions.
Appendix A Robustness Checks
A.1 Adding Soil Covariates in the United States
@@
We found that including soil covariates as controls did not substantially influence the results from the causal forest-based analyses in the United States. Figure A2 shows scatter plots of the results from our causal forest-based analyses when excluding soil covariates from the control vector (as is done in the main text of the paper) versus when including the 5 aforementioned soil covariates in the control vector . As can be seen in the figure, the results do not change much when including soil covariates in the control vector. In addition, in nearly all estimates visualized in Figure A2, controlling for soil covariates did not change whether the 95% confidence interval for that estimate lies above , below , or contains . The only exception is that for the Spring WheatSoybean rotation, the confidence interval for the temperature heterogeneity coefficient ( from Equation (3)) lies above when controlling for soil covariates but contains when not controlling for soil covariates.
A.2 Using a different vegetation index than GCVI
@@
In this appendix we check whether our results were sensitive to the choice of using peak GCVI as a proxy for yield. In particular, we repeated the analyses described in Sections 4.2–4.7, except we used the near infrared reflectance of vegetation (NIRv) as the vegetation index rather than GCVI. NIRv, a vegetation index developed in [60], is the product of the Normalized Difference Vegetation Index and the NIR band, and NIRv has shown strong performance for predicting crop yields [61]. For this senstivity analysis, after removing observations with outlier values for peak harmonic NIRv, we set normalized peak harmonic NIRv to be our outcome variable and ran the causal forest analyses as described in Sections 4.4–4.6. We then converted our results from the normalized NIRv scale to the crop yield scale using the approach described in Section 4.7 (by setting to be the average normalized peak NIRv for crop type in each year and subnational unit when fitting the regression in Equation (4)). We remark that when fitting the calibration model in Equation (4), including with subnational-level yield data in France and Canada (Section H), peak GCVI generally had a higher correlation with the subnational-level crop yield data than did peak NIRv, and thus GCVI-based results are presented in the main text rather than NIRv-based results.
@@
We found that using GCVI as a vegetation index rather than NIRv did not substantially influence the results from the causal forest-based analyses. Figure A3 shows scatter plots of the results from our causal forest-based analyses when using GCVI as the vegetation index (as is done in the main text of the paper) versus when using NIRv as the vegetation index. As can be seen in Figure A3, there is generally good agreement between estimates based on GCVI and those based on NIRv. Although there are some disagreements in the estimates, we remark that these disagreements do not substantially change the main GCVI-based conclusions presented in the paper. In particular, none of our estimates change from being positive and statistically significant to being negative and statistically significant (or vice-versa) when switching from GCVI to NIRv. Additionally, the statistical significance and sign of the precrop effect estimates were the same for all 36 precrop, outcome crop, and country combinations with the exception of the two rotations in Canada with “Pasture and Forages” as the precrop and the Winter WheatCorn rotation in France. In these three cases, the two year sequence was estimated to have a statistically significant, negative effect on yield in the main text, but the estimates were rendered non-statistically significant when using NIRv instead of GCVI.
A.3 Converting outcome variable to the log scale
@@
Because for near , the above quantity on the right hand side is roughly the average percent effect of rotation on the yield estimates (divided by a factor of 100) written as a function of the covariates in . Therefore under the assumption of no unmeasured confounders, can roughly be thought of as the average percent effect of rotation on yields, written as a function of the covariates in . Similar to the analyses in Sections 4.4 and Sections 4.6 we use overlap weighted means of to estimate average rotation benefits and diversification benefits (with different choices of for each estimate). Similar to the analysis in Section 4.5, we assessed heterogeneity with weather on the sample where the propensity scores were between 0.05 and 0.95 by using the best_linear_projection function in the grf package [62] to fit the following modification of the model in equation (3)
We then rescaled the fitted estimates for and using the same approach as in Section 4.5.
Our estimates from these analyses using log estimated yield as the outcome variable were multiplied by a factor of 100 and are plotted in Figure A4 against our estimates from the main text. We find a high agreement between the two approaches meaning that our reported estimates can also be interpreted as relative effects rather than just absolute effects. Notably, this suggests that our finding of higher rotation benefits at higher precipitation levels was not merely driven by greater crop productivity at higher preciptation levels and was instead driven by fundamental increases in the effectiveness of crop rotation in rainier conditions.
A.4 Controlling for geography

### v2

@@ A.5 Analyzing heterogeneity of rotation benefits with weather
The above regression includes a linear time trend to control for gradually changing factors besides weather that may affect rotation benefits (e.g., soil quality, management practices, crop variety choices, etc.), and to reduce how much such factors can confound our estimates of how rotation benefits vary with weather. The regression was fit on the subsample of points where the two year sequence was either AB or BB and the estimated propensity score was between and . Instead of implementing this regression directly, we implemented the regression using the best_linear_projection function in the grf package [38] which provided standard errors and -values that appropriately accounted for the fact that the same sample was being used to estimate the response variable () and to estimate the regression coefficients themselves. The best_linear_projection function also appropriately accounted for the large errors (associated with nonparametric estimation) in the estimates of when computing standard errors and -values of the regression coefficients in (3). In addition, we used the default setting where standard errors and -values for the regression coefficients were based on heteroskedasticity-robust (HC3) estimation, although they did not account for potential spatial correlations (see Section C.5). We refer readers to the documentation of the grf package [38] and [72] for more details on how this regression was fit using an alternative regression that appropriately accounts for errors in estimation of and to [73] for more details on how HC3 standard errors are calculated in regression settings.
@@
We repeated this process for each precrop A, outcome crop B, and country listed in Table S9. The rescaled heterogeneity coefficients can be roughly interpreted as the amount by which the estimated benefit of rotation on our normalized yield proxy are expected to change as the growing season precipitation (or temperature) increases from the 25th percentile to the 75th percentile in the regions where both the rotation AB and the no rotation control BB are commonly observed.
@@
We remark that our approaches do not depend upon linearity assumptions. In spite of our use of a linear model in Equation (3) to study heterogeneity with weather, we do not use linear approaches for the initial estimation of or the treatment effects (e.g., [74]) to ensure that our results on the precrop and diversification effects (Figures 3 and 5; Table S9 (Columns 6 and 7)) do not lean on any linearity assumptions. Moreover, the population regression coefficients giving the best fit to model (3) are still well defined estimands of interest regardless of whether or not is linear in temperature and precipitation. Regardless of whether is linear, the estimands are the well-defined solutions to
@@
and moreover, such estimands still provide useful summary statistics about how varies with precipitation and temperature. We use linear approaches to study heterogeneity in in part because doing so allows valid -value and confidence interval construction. In particular, the best_linear_projection function we use returns confidence intervals that account for the fact that is not known precisely and was estimated from the same sample, whereas existing software for nonlinear approaches (e.g., partial dependence plots, accumulated local effect plots, Shapley values) do not return such confidence intervals.
A.6 Assessment of diverse crop rotations
@@
For each two year crop sequence AB with A and B being distinct, we also assessed whether additional rotational diversity benefits crop yields. To do this, we considered samples with a three year cropping sequence of BAB as simple rotation control units and set on such samples. All samples with a three year cropping sequence of DAB, where D is any croptype that is distinct from A and B were considered as treated units with (we remark that samples of the form DAB where D was fallow, pasture or some other non-crop land classification were not included as treated units and were removed from this analysis (see Appendix I)). With a binary treatment variable distinguishing simple and diverse rotations and an outcome variable giving the GCVI-based proxy (defined in Section A.2) for B’s yield in the third year of the crop sequence, we ran a causal forest and estimated the ATO using the same software and weather controls described in Section A.4. The ATO estimates provided estimates of the benefit of additional rotational diversity compared to a less diverse BAB cropping sequence. While there are many possible diverse rotations with different choices of crops to be grown prior to the AB sequence that can be studied, our ATO estimate gave a weighted average of the diversification benefit across all the crops distinct from A and B that are grown before the AB sequence (with weights proportional to the prevalence in our sample). Finally, as a supplementary analysis, we used the same approach described in Section A.5 to study how the impact of switching from (a simplified sequence of the form BAB) to (a diversified sequence of the form DAB) varied with weather.
@@
A.7 Converting results to the crop yield scale
Our estimated effects of rotation and diversification described in Sections A.4 and A.6 and our rescaled heterogeneity coefficients described in Section A.5 each used normalized peak GCVI as the outcome variable. To improve the interpretability of our results, we converted these estimated effects and heterogeneity coefficients from a unitless GCVI scale to the crop yield scale by using a post-hoc, linear calibration with subnational-level crop yield data.
In particular, to calibrate our results to the yield scale in the United States, for each crop type , county , and year we computed the average normalized peak GCVI, which we call using the sample described in Section A.1. We then extracted annual county-level crop yield data from the National Agricultural Statistics Service of the United States Department of Agriculture [39], which we call . Then for each crop type , we fit the following weighted linear regression to estimate calibration coefficients
@@ A.7 Converting results to the crop yield scale
Due to lack of availability of sufficiently high resolution subnational-level yield data in China, we were unable to deploy the above approach in each country. Instead, we used the scaling factors of learned in the US to convert our results to the yield scale (in units of percent of average yield) in each country for each outcome crop . We used subnational-level yield data in Canada and France to validate this approach and found that subnational-level yield predictions in Canada and France that were based on coefficients from a calibration fit in the US were only slightly less accurate than those based on a calibration fit with local (i.e., from Canada or France) subnational-level yield data (see Appendix J for details).
@@
We remark that we used weighted regression in our calibration to avoid giving too much weight to counties and years that were underrepresented in our US data sample. We also note that we rescale our estimates and standard errors by a factor of rather than by for two reasons. First, the factor that we use converts our estimated average effects (or estimated average changes in effects with weather) to be in units of percent of average yield rather than units of tons per hectare. The former units have been used in recent work studying the impacts of tillage on crop yield [35] and are more interpretable in comparisons across outcome crops. Second, dividing by leads to a more robust calibration in the sense that the calibration learned in the United States using and can more reliably be used in other countries than can a calibration using just and (Appendix J). Finally, we remark that interested readers can convert our results presented in units of percent of average yield to units of tons per hectare using the average yield estimates for our sample presented in Table S10.
A.8 Net effects versus direct effects terminology
@@
Throughout the text, we present estimates of the net effects of cropping sequences on yield and how those net effects vary with weather. The net effects consist of the direct effect of rotation on crop yield as well as the presumably negative indirect effect of crop rotation on yield that is mediated by changes in chemical input use. Throughout the text, in cases where it is important to distinguish between the net effects on yields that we estimate and the direct effects on yields (i.e., the cropping sequence effects at fixed chemical input levels) that are often studied in field experiments, we refer to the latter as “direct”. The net precrop (or diversification) effects we estimate are essentially the precrop (or diversification) effects on yield when the treated units (i.e., the rotated (or diversified) units) are defined to also use their actual reduction in chemical inputs compared to the control units (i.e., the nonrotated (or non-diversified) units). The precrop (and diversification) effects we study are causal quantities of interest, but we note that they only capture the yield benefits of rotation and do not capture additional benefits associated with reduced chemical input costs.
Supplementary Materials: Appendices
Appendix B Discussion of Limitations
The causal inference approach used relies upon the assumption that there are no unmeasured confounders. While we controlled for potential confounders such as weather and geolocation, we did not control for soil quality in our analysis. It is possible that on more fertile land farmers are less likely to rotate their crops, because in such land high crop yields can be obtained without rotation. In Appendix C.1, we find that adding soil covariates from the USDA Soil Survey Geographic Database [75] as controls in the model barely changed our results. This sensitivity check was not conducted in Canada, China, and France due to lack of soil data, but our findings in the United States suggest that omission of soil covariates does not substantially bias our causal effect estimates and our estimates of how they vary with weather.
Nonetheless, our estimates for the benefits of rotation are still useful for understanding the effects of rotating crops on crop yield when using the typical reduction in fertilizer and pesticide use for a given rotation and region. For example, we still find a positive benefit of crop rotation even though less fertilizer and pesticide inputs are typically used in rotated fields. While our estimates do not capture the cost savings from using reduced chemical inputs, they can be coupled with work assessing the impact of rotation on chemical input use and how those impacts vary with weather to obtain a more complete picture of the comprehensive benefits of rotation and how they vary with weather. Furthermore, as shown mathematically in Appendix G, even when there are unmeasured mediators such as fertilizer and pesticide use, our estimates of how the rotation benefits vary with weather could (under certain assumptions) still give unbiased estimates of how the direct rotation benefits vary with weather.
Third, the outcome and treatment variables used do not perfectly capture the actual outcome and treatment variables of interest. The outcome variable is a remotely sensed greenness measure that is merely a proxy for crop yield. In Appendix H, we demonstrate that even though we do not use actual yield estimates as the outcome variable, our approach of using a vegetation index-based outcome variable coupled with an appropriate selection of weather covariates would give the same results, up to an unknown scaling factor, as the observational approach taken in recent works that assess the impacts of crop rotation [12] and tillage [35] on crop yield. Our rescaling of the estimated effects (Section A.7) mitigates the discrepancy between our approach and those taken in [12, 35] by approximating the unknown scaling factor using subnational-level crop yield data. In Appendix H.5, we find that even when using modern, satellite-based yield estimates [59] instead of a vegetation index-based proxy as the outcome variable, the results still roughly differ by an unknown scaling factor, and moreover, for most rotations, county-level regressions of average estimated yield on the average of the proxy gave good approximations of the unknown scaling factors. In addition, in Appendix C.2, we find that our results were robust to the choice of vegetation index that was used as a yield proxy.
Fourth, our estimates for rotation benefits are not perfectly comparable across precrop choices. This is because even within the same country, we only estimated the effect of a rotation on the overlap region where both the treatment and control were commonly observed. For example, when comparing the SoybeanCorn rotation to the Winter WheatCorn rotation in the US, the estimates for the former are for parts of the US where both SoybeanCorn and CornCorn are commonly observed whereas the latter are for parts of the US where both Winter WheatCorn and CornCorn are commonly observed. A similar issue is that the rotation benefits cannot be directly compared across countries even for the same precrop and outcome crop choices. In particular, the analyses in each country corresponded to a different time period based on crop type map availability (Table S1), so observed differences in the rotation benefits between two countries may be partially explained by climatic differences between their study periods. Moreover, the analyses in different countries used crop type maps with different accuracies. For example, we expect rotation benefit estimates from France to suffer from the least attenuation bias compared to those in other countries due to the higher accuracy of crop maps in France (unlike in the US, Canada, and China the crop type maps in France were not based on remote sensing predictions).
Appendix C Robustness Checks
C.1 Adding Soil Covariates in the United States
@@
We found that including soil covariates as controls did not substantially influence the results from the causal forest-based analyses in the United States. Figure S3 shows scatter plots of the results from our causal forest-based analyses when excluding soil covariates from the control vector (as is done in the main text of the paper) versus when including the 5 aforementioned soil covariates in the control vector . As can be seen in the figure, the results do not change much when including soil covariates in the control vector. In addition, in nearly all estimates visualized in Figure S3, controlling for soil covariates did not change whether the 95% confidence interval for that estimate lies above , below , or contains . The only exception is that for the Spring WheatSoybean rotation, the confidence interval for the temperature heterogeneity coefficient ( from Equation (3)) lies above when controlling for soil covariates but contains when not controlling for soil covariates.
C.2 Using a different vegetation index than GCVI
@@
In this appendix we check whether our results were sensitive to the choice of using peak GCVI as a proxy for yield. In particular, we repeated the analyses described in Sections A.2–A.7, except we used the near infrared reflectance of vegetation (NIRv) as the vegetation index rather than GCVI. NIRv, a vegetation index developed in [67], is the product of the Normalized Difference Vegetation Index and the NIR band, and NIRv has shown strong performance for predicting crop yields [68]. For this senstivity analysis, after removing observations with outlier values for peak harmonic NIRv, we set normalized peak harmonic NIRv to be our outcome variable and ran the causal forest analyses as described in Sections A.4–A.6. We then converted our results from the normalized NIRv scale to the crop yield scale using the approach described in Section A.7 (by setting to be the average normalized peak NIRv for crop type in each year and subnational unit when fitting the regression in Equation (4)). We remark that when fitting the calibration model in Equation (4), including with subnational-level yield data in France and Canada (Section J), peak GCVI generally had a higher correlation with the subnational-level crop yield data than did peak NIRv, and thus GCVI-based results are presented in the main text rather than NIRv-based results.
@@
We found that using GCVI as a vegetation index rather than NIRv did not substantially influence the results from the causal forest-based analyses. Figure S4 shows scatter plots of the results from our causal forest-based analyses when using GCVI as the vegetation index (as is done in the main text of the paper) versus when using NIRv as the vegetation index. As can be seen in Figure S4, there is generally good agreement between estimates based on GCVI and those based on NIRv. Although there are some disagreements in the estimates, we remark that these disagreements do not substantially change the main GCVI-based conclusions presented in the paper. In particular, none of our estimates change from being positive and statistically significant to being negative and statistically significant (or vice-versa) when switching from GCVI to NIRv. Additionally, the statistical significance and sign of the precrop effect estimates were the same for all 36 precrop, outcome crop, and country combinations with the exception of the two rotations in Canada with “Pasture and Forages” as the precrop and the Winter WheatCorn rotation in France. In these three cases, the two year sequence was estimated to have a statistically significant, negative effect on yield in the main text, but the estimates were rendered non-statistically significant when using NIRv instead of GCVI.
C.3 Converting outcome variable to the log scale
@@
Because for near , the above quantity on the right hand side is roughly the average percent effect of rotation on the yield estimates (divided by a factor of 100) written as a function of the covariates in . Therefore under the assumption of no unmeasured confounders, can roughly be thought of as the average percent effect of rotation on yields, written as a function of the covariates in . Similar to the analyses in Sections A.4 and Sections A.6 we use overlap weighted means of to estimate average rotation benefits and diversification benefits (with different choices of for each estimate). Similar to the analysis in Section A.5, we assessed heterogeneity with weather on the sample where the propensity scores were between 0.05 and 0.95 by using the best_linear_projection function in the grf package [38] to fit the following modification of the model in equation (3)
We then rescaled the fitted estimates for and using the same approach as in Section A.5.
Our estimates from these analyses using log estimated yield as the outcome variable were multiplied by a factor of 100 and are plotted in Figure S5 against our estimates from the main text. We find a high agreement between the two approaches meaning that our reported estimates can also be interpreted as relative effects rather than just absolute effects. Notably, this suggests that our finding of higher rotation benefits at higher precipitation levels was not merely driven by greater crop productivity at higher preciptation levels and was instead driven by fundamental increases in the effectiveness of crop rotation in rainier conditions.
C.4 Controlling for geography